Waterways Problem 03

Because most of the parts for its irrigation systems are standard, Waterways handles the majority of its manufacturing as a process cost system. There are multiple process departments. Three of these departments are the Molding, Cutting, and Welding departments. All items eventually end up in the Packaging Department, which prepares items for sale in kits or individually.

The following information is available for the Molding department for January.
Work in process beginning:
     Units in process 23,000
     Stage of completion for materials 80%
     Stage of completion for labor and overhead 30%
     Costs in work in process inventory:
       Materials $169,100
       Labor 67,890
       Overhead 17,360
Total costs in beginning work in process $254,350
Units started into production in January 60,000
Units completed and transferred in January 57,500
Costs added to production:
     Materials $263,640
     Labor 292,815
     Overhead 60,300
Total costs added into production in January $616,755
Work in process ending:
     Units in process 25,500
     Stage of completion for materials 50%
     Stage of completion for labor and overhead 10%
Show the equivalent units for materials and conversion costs if Waterways used FIFO instead of weighted-average.
Equivalent Units
Physical Units Materials Conversion
Units accounted for
Completed and transferred out
     Work in process, Jan. 1
     Started and completed in Jan.
Work in process, Jan. 31
     Total units
Answer #1

Show the equivalent units for materials and conversion costs if Waterways used FIFO instead of weighted-average.

Equivalent Units
Physical Units Materials Conversion
Units accounted for
Completed and transferred out
     Work in process, Jan. 1 23000 (23000*20%)= 4600 (23000*70%)= 16100
     Started and completed in Jan. 34500 34500 34500
Work in process, Jan. 31 25500 (25500*50%)= 12750 (25500*10%)= 2550
     Total units 83000 51850 53150

Units started and completed in Jan.= Units completed and transferred-Work in process, Jan 1

= 57500-23000= 34500 units